OpenStudy (anonymous):

Why is the function f(x)=(1/2)x discontinuous

OpenStudy (anonymous):

no. the answer says its discontinuous at even integers but the graph is just a straight line so i dont understand

OpenStudy (aum):

\[f(x) = \frac 12 x\] is the equation of a straight line. It is continuous everywhere. The answer is wrong.

OpenStudy (aum):

Can you post a screenshot of the question and where it says "discontinuous at even integers" ?

OpenStudy (anonymous):

it claims to not be continuous at even numbers because: lim (1/2)x=n x->2n+ but lim (1/2)x=n-1 x->2n-

OpenStudy (aum):

Can you post a screenshot of the question. I think this is NOT f(x) = 1/2 * x. You are omitting some information and a screenshot will help see what is being omitted in your question.

OpenStudy (aum):

See, I knew it!!!

OpenStudy (aum):

That is the greatest integer function as indicated by the square brackets around [1/2x].

OpenStudy (anonymous):

I've never seen this before..

OpenStudy (anonymous):

Ok now i understand it. Thank you so much

OpenStudy (aum):

The greatest integer function f(x) = [x] is the greatest integer less than or equal to x.

OpenStudy (aum):

f(x) = [x] If x is an integer, then f(x) = x If x is not an integer, then f(x) = the nearest integer to the LEFT of it.

OpenStudy (aum):

[5] = 5 [149] = 149 [10.3] = 10 [-15.1] = -16 [5/4] = 1 [-5/4] = -2